# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ A.7. PSLQ integer relation search
This subsection and the next complete the proof of Proposition 6.3. The 70-digit zero ordinates of Appendix B, certified by ARB (Appendix E), serve as input to the PSLQ algorithm.
Integer relation detection uses the PSLQ algorithm (Ferguson–Bailey–Arno [3]) via mpmath.pslq. The 70-digit ARB-certified zero ordinates from Appendix B serve as input at 80-digit working precision.
@@
Remark A.3 (Cautionary experiment).
As a precision control, PSLQ was also applied to the original $20$-digit values at $40$-digit working precision with $\max|n_{k}|\leq 1000$. This returned the vector
[…]
with $\max|n_{k}|=19$ and $\ell^{1}$ norm $184$, whose weighted sum vanishes to all $20$ displayed digits. Evaluating at $70$-digit precision gives $\sum n_{k}\gamma_{k}^{\prime}=2.72\times 10^{-18}$, unambiguously nonzero. The spurious detection is explained by the effective precision of a linear combination of $20$ terms with $\ell^{1}$ norm $184$: $20-\log_{10}(184)\approx 17.7$ digits, below the residual threshold. The $70$-digit certification provides ample margin above this threshold.
@@ A.8. Bessel tail bound
For any integer vector $\mathbf{n}$ with $\max|n_{k}|\geq N=1001$, the standard bound $|J_{N}(x)|\leq(x/2)^{N}/N!$ gives, at the most pessimistic integration range $T=2000$ and with $b_{\max}=b_{1}=4.499\times 10^{-2}$:
$\left(\frac{b_{1}\cdot 2000}{2}\right)^{1001}\frac{1}{1001!}<10^{-916}.$
The count of integer vectors with $\max|n_{k}|=N$ in dimension $20$ is at most $(2N+1)^{20}\leq(2003)^{20}<10^{66}$. Summing over all $N\geq 1001$ gives total resonance contribution bounded by $10^{66}\cdot 10^{-916}=10^{-850}$.
The resonance correction is bounded by $10^{-850}$ and does not affect any reported digit of $f_{S_{L}^{(20)}}(0)=8.3129$. The PSLQ search (§A.7) and Bessel tail bound together certify $d_{20}=0$. This completes the proof of Proposition 6.3. ∎

### 12 Mar 2026 v3

@@ A.7. PSLQ integer relation search
This subsection and the next complete the proof of Proposition 6.3. Integer relation detection uses the PSLQ algorithm (Ferguson–Bailey–Arno [3]) via mpmath.pslq. The 70-digit ARB-certified zero ordinates from Appendix B serve as input at 80-digit working precision.
@@ A.8. Bessel tail bound
For any integer vector $\mathbf{n}$ with $\max|n_{k}|\geq N=1001$, the standard bound $|J_{N}(x)|\leq(x/2)^{N}/N!$ gives, at the most pessimistic integration range $T=2000$ and with $b_{\max}=b_{1}=4.499\times 10^{-2}$ (upper endpoint of the ARB ball):
$\left(\frac{b_{1}\cdot 2000}{2}\right)^{1001}\frac{1}{1001!}<10^{-915.8}.$
The sum over all $N\geq 1001$ acquires a geometric series factor $1/(1-b_{1}\cdot 1000/1002)<1.05$. The count of integer vectors with $\max|n_{k}|=N$ in dimension $20$ is at most $(2N+1)^{20}\leq(2003)^{20}<10^{66.1}$. Summing over all $N\geq 1001$ gives total resonance contribution bounded by $10^{66.1}\cdot 10^{-915.8}\cdot 1.05<10^{-849.5}$, ARB-certified with $\log_{10}(\text{total})\in[-849.759\pm 10^{-72}]$.
The resonance correction is bounded by $10^{-849.5}$ and does not affect any reported digit of $f_{S_{L}^{(20)}}(0)=8.3129$. The PSLQ search (§A.7) and Bessel tail bound together certify $d_{20}=0$. This completes the proof of Proposition 6.3. ∎
@@ A.9. Reproducibility
Ancillary files. All ancillary Python scripts, including self-contained ARB certificates for individual theorems, are deposited at https://doi.org/10.5281/zenodo.18783098 and mirrored at https://github.com/PeterShiller/Persistent_Heuristics_Part_I. Each script requires only python-flint $\geq 0.8.0$ and sympy $\geq 1.12$; no external data files are needed.